1. Suppose Leonard, Nixon, & Shull Corporation’s projected free cash flow for next year is $100,000, and FCF is expected to grow at a constant rate of 6%. If the company’s weighted average cost of capital is 11%, what is the value of its operations?

a. $1,714,750
b. $1,805,000
c. $1,900,000
d. $2,000,000
e. $2,100,000
1
Expert's answer
2012-02-16T09:27:07-0500
a. $1,714,750
b. $1,805,000
c. $1,900,000
d. $2,000,000
e. $2,100,000


Solution:

Calculate Horizon value = FCF*(1+ growth rate)/(WACC – growth rate) = $100 000 * 1,06 / (0,11 – 0,6) =
=$2 120 000

Calculate Present value = CF / (1+WACC)^t, t – time period

[table]Yearly earnings Present value100 000,0090 090,092 120 000,001 909 909,91[/table]
The value of operation = ∑ CFs = $90 090,09 + $1 909 909,91 = $2 000 000